OpenStudy (anonymous):

01.04 Graphing Linear Equations and Inequalities

OpenStudy (anonymous):

Need any help sir?

OpenStudy (anonymous):

ya i think i do

OpenStudy (anonymous):

Alright, what's up?

OpenStudy (anonymous):

-x - 2y = 6 how do i graph this?

OpenStudy (anonymous):

Ok, first let's put it into a form that we can recognize the slope and y intercept better. Slope-Intercept form: y=mx+b

OpenStudy (anonymous):

-2y=x+6

OpenStudy (anonymous):

If we add x to both sides, we get -2y = 6 + x, rearrange the x and 6 to x + 6, so we have -2y = x + 6 That's looking a bit better. Do you know how to get y by itself on the left side after that?

OpenStudy (anonymous):

Ah, great! Let's get y to be by itself now.

OpenStudy (anonymous):

What would you do?

OpenStudy (anonymous):

divide by -2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Awesome, so we get... y = -(x + 6)/2

OpenStudy (anonymous):

Sorry, I forgot the (-).

OpenStudy (anonymous):

so next i would have to find slope?

OpenStudy (anonymous):

So -(x+6)/2 is the same as -x/2 + 6/2 correct?

OpenStudy (anonymous):